# Graph sub-sampling for divide-and-conquer algorithms in large networks

### v2

@@ 5.2 Core-periphery detection
We now compare the sub-sampling routines using the CP divide-and-conquer algorithm on real-world networks: Airport (Csardi,, 2013) is a network where nodes correspond to airports and edges to flights between airports ($n=755$, $m=4623$); and Twitch (Rozemberczki et al.,, 2019; Leskovec and Krevl,, 2014) has nodes which are users and edges represent friendships ($n=168,11$, $m=6,797,557$). Edge weights and directions were again removed, along with self-loops.
For each network, we report the size of the core and corresponding BE metric value for the optimal core returned by the algorithm. In the simulation settings, we simply used the core proportions as a measure of a node’s coreness, but for the real data, we want binary CP labels. To convert the core proportions to binary labels, we sort the nodes in descending order of their proportions. Then one by one, starting with the node with the largest proportion, we add this node to the core and compute the BE metric. We continue this process and select the core corresponding to the largest BE metric.
While the airport network is relatively small, this allows us to also find the CP labels using the vanilla BE metric algorithm without the divide-and-conquer step. We label this as “Full” in our results, and this allows us to compare both the core nodes and BE metric returned by the divide-and-conquer algorithm. Finally, we note the following values of $q$ and $B$ that we used for each network: Airport, $q=100/n$ and $B=5000$; and Twitch, $q=500/n$ and $B=25,000$.
In Table 5, we report the core size and BE metric for the detected cores, and in Figure 3, we compare the cores returned by the different algorithms. Specifically, for each sub-sampling pair, we compute the Jacaard coefficient (JC) between the core sets. The more similar the cores are, the larger the JC will be. For Airport, RE returns the most similar core as that of the full algorithm. All other methods have similar core sizes and BE metric values. Indeed, Figure 3 shows that each core is not only a comparable size, but the core nodes returned by each sub-sampling method have significant overlap. For Twitch, RE and RW yield the largest BE values, with RN, BFS and RNN also being similar. DN yields a vastly different core size and BE value, while DFS does not perform as well. For this network, RW and RE have the most similar cores.
@@
Full 33 0.236
RN 29 0.236 275 0.077
DN 31 0.233 164 0.004
RE 31 0.233 170 0.079
BFS 35 0.234 88 0.076
DFS 27 0.225 155 0.066
RNN 30 0.234 91 0.076
RW 32 0.235 141 0.079
Table 5: Real data analysis results for core-periphery divide and conquer algorithm. $k$ is the number of nodes assigned to the core and BE is the value of the Borgatti and Everett metric corresponding to the optimal labels.
@@ 6 Conclusion
Based on our findings, we recommend random node sampling when trying to detect community structure. This method had some of the best performance in the simulated networks, as well as in the real-world networks. Moreover, random node sampling was studied theoretically in Mukherjee et al., (2021) and is the easiest sub-sampling method to implement. The reason for this good performance could be because random node sampling can easily sample across the network without getting “stuck” in any one part. Of course, random node sampling, as well as the other methods, will suffer if there are communities with only a few nodes. In this case, it is important to increase the size of the sub-graphs and/or the number of iterations. Randomly sampling nodes proportional to their degree as well as random node neighbor sampling also provided good results on the synthetic and real-world networks. Breadth-first and depth-first search algorithms, however, should be avoided for this task.
As for core-periphery identification, there was no clear winner, but samplers which selected core nodes with a higher probability consistently performed the best. Random edge sampling and random walk both yielded good results across the simulated and real-world data. On the other hand, random node sampling and random node neighbor performed quite poorly in these settings, aligning with our theoretical results. Finally, the real-data results show that many different cores can yield similar objective function values, implying the objective function surface is relatively “flat.”
The contrasting performance of sub-sampling routines on the different identification task is revealing. Random node and random node neighbor sampling performed the best for the community detection problem, but were the two weakest methods for core-periphery identification. On the other hand, breadth-first and depth-first searches are good approaches for identifying core-periphery structure, but not for community structure. These results highlight the importance of tailoring the choice of sub-sampling algorithm to the particular problem. Indeed, our CP theory revealed the optimal feature for a sampler is a high probability of sampling core nodes. Deriving similar theory for future problems will help researchers choose the best methods, or even may even drive them to derive novel sub-sampling schemes for the specific problem.
@@
Proof of Theorem 3.2:
Let $\boldsymbol{c}^{*}$ be the true CP labels. Let $\hat{\boldsymbol{c}}$ be the estimated labels from divide-and-conquer procedure where
@@
By Cauchy-Schwarz, we then have
@@
is the mis-classification error of the algorithm on a single sub-graph. We can get an error here from one of two ways. First, if our sub-graph contains node $i$ but incorrectly classify it. Second, if we don’t sample node $i$ and it’s in the core. Thus,
$\frac{1}{n}\sum_{i=1}^{n}\{\hat{c}_{i}^{(b)}-c_{i}^{*}\}^{2}=\frac{1}{n}||\hat{\boldsymbol{c}}^{(b)}-\boldsymbol{c}^{(b)*}||_{2}^{2}+\frac{1}{n}\sum_{i=1}^{n}y_{i}^{(b)}$
where $\boldsymbol{c}^{(b)*}$ is the true core labels restricted to sub-graph $b$ and $y_{i}^{(b)}=1$ if node $i$ is in the core and was not sampled in $b$, and 0 otherwise. Thus,
$\delta(\hat{\boldsymbol{c}},\boldsymbol{c}^{*})\leq\frac{1}{nB}\sum_{b=1}^{B}\left(||\hat{\boldsymbol{c}}^{(b)}-\boldsymbol{c}^{(b)*}||_{2}^{2}+\sum_{i=1}^{n}y_{i}^{(b)}\right)$
@@
$\mathbb{E}\delta(\hat{\boldsymbol{c}},\boldsymbol{c}^{*})\leq\frac{1}{n}(\mathbb{E}||\hat{\boldsymbol{c}}^{(S)}-\boldsymbol{c}^{(S)*}||_{2}^{2}+\mathbb{E}\sum_{i=1}^{n}y_{i}^{(S)})$ (8)
where $S$ is a randomly chosen sub-graph from our sub-sampling routine. This gives us the desired result.
Proof of Corollary 3.2:
Here, we present the specific error for each of the sub-sampling routines. First, let $Y^{(b)}$ be the number of core nodes drawn in sub-sample $b$, i.e.,
$Y^{(b)}=k-\sum_{i=1}^{n}y_{i}^{(b)}.$
Without loss of generality, assume that the first $k$ nodes in the network are the core nodes.
Random Node: For the random node sampling scheme, each node is equally likely to be selected. Thus, $Y^{(b)}$ has a hyper-geometric distribution with a total population of size $n$,“success” states in the population $k$, and draws $qn$, so
$\mathbb{E}Y^{(b)}=\frac{qnk}{n}=qk.$
Another way to see this is that the probability of a core node being drawn is $k/n$ and we draw $qn$ nodes so we get the $qk$. Thus, the expected number of core nodes which are not sampled is simply
@@
$\mathbb{E}\ \frac{1}{n}\sum_{i=1}^{n}y_{i}^{(b)}=(1-q)\frac{k}{n}.\ \square$
@@
For the following derivations, we note that $m=\sum_{i<j}A_{ij}$ is highly concentrated around its mean, $\sum_{i<j}P_{ij}$, and the error bound for $m$ is much smaller than that of the other (appropriately scaled) random variables. Therefore, in what follows, we will ignore the randomness of $m$ and consider it to be approximately equal to its expectation, i.e., if the data comes from a CP-SBM, then
$2m\approx\mathbb{E}2m=k(k-1)p_{11}+2k(n-k)p_{12}+(n-k)(n-k-1)p_{22}.$
We also assume that the network is large enough such that the probability of sampling the same node twice is negligible.
Degree Node: For Random Node, each node had a $1/n$ probability of being selected. Now, a nodes’ the probability of being drawn is proportional to its degree. We can model this as drawing from an urn with $2m$ balls where there are $d_{i}$ balls labeled for node $i$, where $d_{i}$ is the degree of node $i$. Then $Y^{(b)}$ has a hyper-geometric distribution where now there are $2m$ population elements, $\sum_{i=1}^{k}d_{i}$ success states, and $qn$ draws. Thus,
$\mathbb{E}Y^{(b)}=\mathbb{E}\mathbb{E}(Y^{(b)}|d_{i})=\mathbb{E}\frac{qn\sum_{i=1}^{k}d_{i}}{2m}=\frac{qnk\{(k-1)p_{11}+(n-k)p_{12}\}}{k(k-1)p_{11}+2k(n-k)p_{12}+(n-k)(n-k-1)p_{22}}$
since
$\mathbb{E}d_{i}=(k-1)p_{11}+(n-k)p_{12}$
if node $i$ is in the core. Thus,
$\mathbb{E}\ \frac{1}{n}\sum_{i=1}^{n}y_{i}^{(b)}=\left[1-\frac{qn\{(k-1)p_{11}+(n-k)p_{12}\}}{k(k-1)p_{11}+2k(n-k)p_{12}+(n-k)(n-k-1)p_{22}}\right]\frac{k}{n}$
For large $n$, this simplifies to
$\left(1-q\frac{p_{12}}{p_{22}}\right)\alpha_{n}.\ \square$
Random Edge: We can again imagine an urn filled with balls where now the balls are labeled with edges instead of nodes. For each ball that we draw, we sample the two nodes which connect this edge, so we only make $qn/2$ draws. There are two types of edges that we are interested in: a core-core edge where both nodes are in the core, and a core-periphery edge, where one node is from the core. If $M_{cc}^{(b)}$ and $M_{cp}^{(b)}$ are the number of core-core and core-periphery edges drawn in sub-sample $b$, respectively, then
$\mathbb{E}Y^{(b)}=2\mathbb{E}M_{cc}^{(b)}+\mathbb{E}M_{cp}^{(b)}.$
Let $m_{cc}$ and $m_{cp}$ be the observed number of core-core and core-periphery edges in the network. Then $M_{cc}^{(b)}$ has a hyper-geometric distribution with population size $m$, $m_{cc}$ success states and $qn/2$ draws. Similarly, $M_{cp}^{(b)}$ has a hyper-geometric distribution with population size $m$, $m_{cp}$ success states and $qn/2$ draws. Thus,
$\mathbb{E}Y^{(b)}=2\mathbb{E}\mathbb{E}(M_{cc}^{(b)}|m_{cc})+\mathbb{E}\mathbb{E}(M_{cp}^{(b)}|m_{cp})\\ =2\mathbb{E}\frac{m_{cc}qn}{2m}+\mathbb{E}\frac{m_{cp}qn}{2m}=\frac{k(k-1)p_{11}qn}{2m}+\frac{k(n-k)p_{12}qn}{2m}$
Therefore,
$\mathbb{E}\ \frac{1}{n}\sum_{i=1}^{n}y_{i}^{(b)}=\left[1-\frac{qn\{(k-1)p_{11}+(n-k)p_{12}\}}{k(k-1)p_{11}+2k(n-k)p_{12}+(n-k)(n-k-1)p_{22}}\right]\frac{k}{n}$
which is the same result as in Degree Node. Again, for large $n$, this simplifies to
$\left(1-q\frac{p_{12}}{p_{22}}\right)\alpha_{n}.\ \square$
Random Node Neighbor For Random Node Neighbor, we randomly sample a node and then include that node as well as all of its neighbors in the sub-graph. First, let’s find the expected number of core nodes per draw. The expected number of core nodes drawn for a single iteration of this sampler depends on whether the initial node as in the core or periphery. Thus, assuming an SBM,
$\mathbb{E}(\text{core nodes}|\text{core node})\mathbb{P}(\text{core node})+\mathbb{E}(\text{core nodes}|\text{periphery node})\mathbb{P}(\text{periphery node})$
=$=\{1+(k-1)p_{11}\}\frac{k}{n}+kp_{12}\frac{n-k}{n}$
$=C_{n}$
So for each draw, we expect to include $C_{n}$ core nodes in the sub-graph. But how many draws do we make? For each draw, the expected number of nodes sampled again depends on the first draw.
$\mathbb{E}(\text{nodes}|\text{core node})\mathbb{P}(\text{core node})+\mathbb{E}(\text{nodes}|\text{periphery node})\mathbb{P}(\text{periphery node})$
$=1+\{(k-1)p_{11}+(n-k)p_{12}\}\frac{k}{n}+\{kp_{12}+(n-k-1)p_{22}\}\frac{n-k}{n}$
$=T_{n}$
where the first 1 is from the node we initially drew. Thus, if we want to sample $qn$ nodes, then we expect to make $qn/T_{n}$ draws. And for each draw, we expect to sample $C_{n}$ core nodes. Thus, the expected number of core nodes in our sub-graph is
$qn\frac{C_{n}}{T_{n}}=qn\frac{\{1+(k-1)p_{11}\}\frac{k}{n}+kp_{12}\frac{n-k}{n}}{1+\{(k-1)p_{11}+(n-k)p_{12}\}\frac{k}{n}+\{kp_{12}+(n-k-1)p_{22}\}\frac{n-k}{n}}$
Therefore,
$\mathbb{E}\ \frac{1}{n}\sum_{i=1}^{n}y_{i}^{(b)}=\left[1-qn\frac{\{1+(k-1)p_{11}\}\frac{1}{n}+p_{12}\frac{n-k}{n}}{1+\{(k-1)p_{11}+(n-k)p_{12}\}\frac{k}{n}+\{kp_{12}+(n-k-1)p_{22}\}\frac{n-k}{n}}\right]\frac{k}{n}.$
For large $n$, this simplifies to
$\mathbb{E}\ \frac{1}{n}\sum_{i=1}^{n}y_{i}^{(b)}=\left(1-q\frac{p_{12}}{\alpha_{n}p_{12}+p_{22}}\right)\alpha_{n}.\ \square$
Random Walk While the calculations are much more involved, we can also find the expected number of core nodes that aren’t sampled for Random Walk. Let $E^{l}$ be the expected number of core nodes traversed on a random walk through a CP-SBM. Then, depending on whether the first node selected is a core or periphery node we have
$E_{l}=\frac{k}{n}(1+x_{l-1})+\frac{n-k}{n}y_{l-1}$
where $x_{l}$ and $y_{l}$ are the expected number of core nodes traversed if the starting node is in the core and periphery, respectively. We can set up a coupled recurrence relationship to solve for these terms. Again, the key observation is that the expectation just depends on whether the next step is to a core or peripheral node:
$x_{l}$ $=\underbrace{\frac{(k-1)p_{11}}{(k-1)p_{11}+(n-k)p_{12}}}_{\alpha}(1+x_{l-1})+\underbrace{\frac{(n-k)p_{12}}{(k-1)p_{11}+(n-k)p_{12}}}_{\beta}y_{l-1}$
$y_{l}$ $=\underbrace{\frac{kp_{12}}{kp_{12}+(n-k-1)p_{22}}}_{\gamma}(1+x_{l-1})+\underbrace{\frac{(n-k-1)p_{22}}{kp_{12}+(n-k-1)p_{22}}}_{\delta}y_{l-1}$
with initial conditions
$x_{1}=1\text{ and }y_{1}=0.$
So we can simplify this as
$x_{l}$ $=\alpha x_{l-1}+\beta y_{l-1}+\alpha$
$y_{l}$ $=\gamma x_{l-1}+\delta y_{l-1}+\gamma$
This can be written in matrix notation as
${\bf v}_{l}={\bf A}{\bf v}_{l-1}+{\bf c}$
where ${\bf v}_{l}=(x_{l},y_{l})^{T}$, ${\bf c}=(\alpha,\gamma)^{T}$ and
${\bf A}=\begin{pmatrix}\alpha&\beta\\ \gamma&\delta\end{pmatrix}$
where ${\bf v}_{1}=(1,0)^{T}$. This has a general solution of
${\bf v}_{l}={\bf A}^{l}{\bf v}_{1}+\sum_{i=1}^{l-1}{\bf A}^{i}{\bf c}.$
Then we just need to plug these solutions into our equation for $E_{l}$, noting that we have $l=qn$ for the length of our random walk.

### v3

@@ 5.2 Core-periphery detection
[…]
For each network, we report the size of the core, $k$, and corresponding Borgatti and Everett (BE) metric value for the optimal core returned by the algorithm, computed using (2). The BE values quantifies the strength of core-periphery structure detected by the algorithm, with larger values implying a better core was found. In the simulation settings, we simply used the core proportions as a measure of a node’s coreness, but for the real data, we desire binary CP labels. To convert the core proportions to binary labels, we sort the nodes in descending order of their proportions. Then one by one, starting with the node with the largest proportion, we add this node to the core and compute the BE metric. We continue this process and select the core corresponding to the largest BE metric.
[…]
In Table 4, we report the core size and BE metric for the detected cores, and in Figure 5, we compare the cores returned by the different algorithms. Specifically, for each sub-sampling pair, we compute the Jacaard coefficient (JC) between the returned core sets. If $S_{1}$ and $S_{2}$ are two sets, then the JC is computed as
$JC(S_{1},S_{2})=\frac{|S_{1}\cap S_{2}|}{|S_{1}\cup S_{2}|}.$
In words, the JC is the proportion of total elements in the two sets which are shared. It takes values from 0 to 1 with 1 indicating that $S_{1}$ and $S_{2}$ are identical. In this way, the more similar the cores are, the larger the JC will be.
For Airport, all method return similar core sizes, cores and BE metric values as that of Full. Moreover, the cores for DN and RE are exactly the same (JC of 1), supporting the results from Corollary 3.2. For Twitch, RE and RW yield the largest BE values, with RN, BFS and RNN also having similar values. DN yields a vastly different core size and BE value, while DFS does not perform as well. For this network, RW and RE have the most similar cores, which supports Lovász, (1993). On the other hand, DN and RE’s results are quite dissimilar, implying that the SBM assumption of Corollary 3.2 may be necessary for the results to hold.
@@
Full 33 0.24
RN 29 0.24 275 0.08
DN 31 0.23 164 0.00
RE 31 0.23 170 0.08
BFS 35 0.23 88 0.08
DFS 27 0.23 155 0.08
RNN 30 0.23 91 0.08
RW 32 0.24 141 0.08
Table 4: Real data analysis results for core-periphery divide and conquer algorithm. $k$ refers to the number of nodes assigned to the core by the respective algorithms, and BE is the value of the Borgatti and Everett metric (2) corresponding to the optimal labels.
@@ 6 Conclusion
Based on our findings, we recommend random node sampling when trying to detect community structure. This method had some of the best performance in the simulated networks as well as on the real-world networks. Moreover, random node sampling was studied theoretically in Mukherjee et al., (2021) and is the easiest sub-sampling method to implement. The reason for this good performance could be because random node sampling can easily sample across the network without getting “stuck” in any one part. Of course, random node sampling, as well as the other methods, will suffer if there are communities with only a few nodes. In this case, it is important to increase the size of the sub-graphs and/or the number of iterations. Randomly sampling nodes proportional to their degree as well as random node neighbor sampling also provided good results on the synthetic and real-world networks. Breadth-first and depth-first search algorithms, however, should be avoided for this task. Additionally our results showed that in many cases, the divide-and-conquer algorithm is not needed at all, as the base algorithm can yield good detection results in a short amount of time.
As for core-periphery identification, there was no clear winner, but samplers which selected core nodes with a higher probability consistently performed the best. Random edge sampling and random walk both yielded good results across the simulated and real-world data. On the other hand, random node sampling and random node neighbor performed quite poorly in these settings, aligning with our theoretical results. The real-data results show that many different cores can yield similar objective function values, implying the objective function surface is relatively “flat.” For CP identification, our results imply that the divide-and-conquer algorithm should always be preferred to the base algorithm as it is faster and is better at identifying the core.
It is important to note that these results were based on our simulated and real-world data experiments and may not generalize to other settings. That being said, the contrasting performance of sub-sampling routines on the different identification task is revealing. Random node and random node neighbor sampling performed the best for the community detection problem, but were the two weakest methods for core-periphery identification. On the other hand, breadth-first and depth-first searches are good approaches for identifying core-periphery structure, but not for community structure. These results highlight the importance of tailoring the choice of sub-sampling algorithm to the particular problem. Indeed, our CP theory revealed the optimal feature for a sampler is a high probability of sampling core nodes. Deriving similar theory for future problems will help researchers choose the best methods, or even may even drive them to derive novel sub-sampling schemes for the specific problem.
Statements and Declarations
The author declares no competing interests nor have any funding to report.
@@
Theorem 3.2:
Given some graph $G$ with adjacency matrix ${\bf A}$ and sampling scheme $\mathcal{S}$, let $S$ be a randomly chosen sub-graph. Let $\hat{\boldsymbol{c}}$ be estimated CP labels using the CP divide-and-conquer algorithm. Then the expected mis-clustering rate, $\mathbb{E}\delta(\hat{\boldsymbol{c}},\boldsymbol{c}^{*})$ can be bounded by
where $\hat{\boldsymbol{c}}^{(S)},(\boldsymbol{c}^{*})^{(S)}\in\{0,1\}^{qn}$ are the estimated and true CP labels, respectively, restricted to sub-graph $S$, and $y_{i}^{(S)}=1$ if node $i$ is in the core and was not sampled in sub-graph $S$, and 0 otherwise.
Proof. Let $\boldsymbol{c}^{*}$ be the true CP labels. Let $\hat{\boldsymbol{c}}$ be the estimated labels from the divide-and-conquer procedure where
@@
By Cauchy-Schwarz (or Jensen’s) inequality, we have
@@
is the mis-classification error of the algorithm on a single sub-graph. We can get an error here from one of two ways: if our sub-graph contains node $i$ but incorrectly classifies it, or if we do not sample node $i$ and it is in the core. Thus,
$\frac{1}{n}\sum_{i=1}^{n}\{\hat{c}_{i}^{(b)}-c_{i}^{*}\}^{2}=\frac{1}{n}||\hat{\boldsymbol{c}}^{(b)}-\boldsymbol{c}^{*(b)}||_{2}^{2}+\frac{1}{n}\sum_{i=1}^{n}y_{i}^{(b)}$
where $\boldsymbol{c}^{*(b)}$ is the true core labels restricted to sub-graph $b$ and $y_{i}^{(b)}=1$ if node $i$ is in the core and was not sampled in $b$, and 0 otherwise. Thus,
$\delta(\hat{\boldsymbol{c}},\boldsymbol{c}^{*})\leq\frac{1}{nB}\sum_{b=1}^{B}\left(||\hat{\boldsymbol{c}}^{(b)}-\boldsymbol{c}^{*(b)}||_{2}^{2}+\sum_{i=1}^{n}y_{i}^{(b)}\right)$
@@
$\mathbb{E}\delta(\hat{\boldsymbol{c}},\boldsymbol{c}^{*})\leq\frac{1}{n}(\mathbb{E}||\hat{\boldsymbol{c}}^{(S)}-\boldsymbol{c}^{*(S)}||_{2}^{2}+\mathbb{E}\sum_{i=1}^{n}y_{i}^{(S)})$ (10)
where $S$ is a randomly chosen sub-graph from our sub-sampling routine. This gives us the desired result. $\square$
Corollary 3.2:
Let ${\bf A}$ be generated from a CP-SBM with $n$ nodes and $k$ core nodes. Let $S$ be a randomly chosen sub-graph of ${\bf A}$ of size $q_{n}n$ where $q_{n}\in(0,1)$ may depend on $n$. Furthermore, let $y_{i}^{(S)}$ is the event that node $i$ is a core node and not sampled in $S$. If RN is used to sample $S$ and $q_{n}=O(1)$, then
$\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}y_{i}^{(S)}=(1-q)\frac{k}{n}.$ (11)
If $S$ is sampled via DN or RE and $q_{n}=o(1)$, then
$\lim_{n\to\infty}\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}y_{i}^{(S)}=\left(1-q_{n}\frac{p_{12}}{p_{22}}\right)\alpha_{n}.$ (12)
Proof.
Random Node: Recall that $q_{n}=q=O(1)$ in this setting. Let $Y^{(\mathcal{S})}$ be the number of core nodes drawn in sub-graph $\mathcal{S}$, i.e.,
$Y^{(\mathcal{S})}=k-\sum_{i=1}^{n}y_{i}^{(\mathcal{S})}.$
Since each node is equally likely to be sampled, then $Y^{(\mathcal{S})}$ has a hyper-geometric distribution with a total population of size $n$, $k$ success states in the population, and $qn$ draws, so
$\mathbb{E}Y^{(\mathcal{S})}=\frac{qnk}{n}=qk.$
Thus, the expected number of core nodes which are not sampled is simply
@@
$\mathbb{E}\ \frac{1}{n}\sum_{i=1}^{n}y_{i}^{(\mathcal{S})}=(1-q)\frac{k}{n}.\ \square$
@@
For the following derivations, we note that $m_{n}=\sum_{i<j}A_{ij}$ is highly concentrated around its mean, $\sum_{i<j}P_{ij}$, and the error bound for $m_{n}$ is much smaller than that of the other (appropriately scaled) random variables. Therefore, in what follows, we will ignore the randomness of $m_{n}$ and consider it to be approximately equal to its expectation, i.e., if the data comes from a CP-SBM, then
$2m_{n}\approx\mathbb{E}2m_{n}=k(k-1)p_{11}+2k(n-k)p_{12}+(n-k)(n-k-1)p_{22}\sim n^{2}p_{22}.$
We also assume that $q_{n}=o(1)$, and, without loss of generality, that the first $k$ nodes are the core nodes. Furthermore, let $z_{j}^{(S)}\in\{0,1\}$ be the random variable such that $z_{j}=1$ if draw $j$ for sub-sample $S$ yields a core node, and 0 otherwise.
Degree Node: When we sample the first node, the probability that it is a core node is
$\mathbb{E}z_{1}^{(S)}=\mathbb{P}(z_{1}^{(S)}=1)=\frac{\sum_{i=1}^{k}d_{i}}{2m_{n}}$
where $d_{i}$ is the degree of node $i$. For core node $i$,
$\mathbb{E}d_{i}=(k-1)p_{11}+(n-k)p_{12}\sim np_{12}$
for large $n$. Thus, we can see that
$\mathbb{P}(z_{1}^{(S)}=1)\sim\frac{k}{n}\frac{p_{12}}{p_{22}}.$
Note that for any $p_{12}/p_{22}$, there exists some $N$ such that for all $n\geq N$, $\frac{k}{n}\frac{p_{12}}{p_{22}}<1$, making this a valid probability. Now, since we are sampling without replacement, the probability of drawing a core node changes on the ensuing draws, depending on whether the previous nodes drawn were core or periphery. However, if we can show that this change in probability is small, then we can model the number of core nodes drawn with the binomial distribution. In particular, assume that we have sampled nodes $v_{1},\dots,v_{qn-1}$, and note that for any draw, the expected degree, $d$, of the node which was sampled is
$\mathbb{E}d$
=$=\{(k-1)p_{11}+(n-k)p_{12}\}\frac{k\{(k-1)p_{11}+(n-k)p_{12}\}}{2m_{n}}$
+$+\{kkp_{12}+(n-k-1)p_{22}\}\frac{(n-k)\{kp_{12}+(n-k-1)p_{22}\}}{2m_{n}}$
$\sim np_{22}$
Thus, on the $q_{n}n$th draw,
$\mathbb{P}(z_{q_{n}n}^{(\mathcal{S})}=1)$ $=\frac{\sum_{i=1}^{k}d_{i}-\sum_{j=1}^{q_{n}n-1}\mathbb{I}(1\leq v_{j}\leq k)d_{v_{j}}}{2m_{n}-\sum_{j=1}^{q_{n}n-1}d_{v_{j}}}$
$\sim\frac{nkp_{12}-q_{n}n\times\frac{k}{n}\frac{p_{12}}{p_{22}}\times np_{22}}{n^{2}p_{22}-q_{n}n(np_{22})}$
$\sim\frac{nkp_{12}(1-q_{n})}{n^{2}p_{22}(1-q_{n})}$
$\sim\frac{k}{n}\frac{p_{12}}{p_{22}}$
since we assumed $q_{n}=o(1)$. Thus, we can model, $Y^{(S)}$, the number of core nodes drawn in sub-graph $S$, as
$Y^{(S)}\sim\mathsf{Binomial}(q_{n}n,\tfrac{k}{n}\tfrac{p_{12}}{p_{22}}),$
so $\mathbb{E}Y^{(S)}=q_{n}k\frac{p_{12}}{p_{22}}$ such that
$\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}y_{i}^{(S)}\sim\left(1-q_{n}\frac{p_{12}}{p_{22}}\right)\alpha_{n}$
as we hoped to show. $\square$
Random Edge: We follow a similar pattern for RE. Recall that we are sampling edges now, so the probability that the first edge we sample is incident to a core node is:
$\mathbb{P}(z_{1}^{(S)}=1)=\frac{\sum_{i=1}^{k}d_{i}}{m}\sim\frac{knp_{12}}{\frac{1}{2}n^{2}p_{22}}\sim 2\frac{k}{n}\frac{p_{12}}{p_{22}}$
Note that since we assume $\alpha_{n}=o(1)$, the probability of drawing an edge incident to two core nodes is vanishingly small. Now, notice that since we draw two nodes every time we sample an edge, we only make $\frac{1}{2}q_{n}n$ draws. Then wimilar to RN, for the $\frac{1}{2}q_{n}n$th draw, the probability of drawing an edge with a core node is still
$\mathbb{P}(z_{q_{n}n}^{(\mathcal{S})}=1)\sim 2\frac{k}{n}\frac{p_{12}}{p_{22}}$
Thus, we can model $Y^{(\mathcal{S})}$, the number of core nodes drawn in sub-graph $\mathcal{S}$ as
$Y^{(\mathcal{S})}\sim\mathsf{Binomial}(\tfrac{1}{2}q_{n}n,2\tfrac{k}{n}\tfrac{p_{12}}{p_{22}})$
and the result follows as before. $\square$
Additional simulation results
We replicate the simulation study from the main manuscript, now using the Walktrap community detection algorithm (Pons and Latapy,, 2005) as the base algorithm. We use the same settings as in the main manuscript. The results are in Figures 6 and 7. We can see that the results are very similar to those with the Fast Greedy base algorithm. The main difference is that the base algorithm outperforms $\mathsf{PACE}$ for all settings (except extremely disproportionate community sizes), but is slower for many settings. These results are more in line with what we would expect: the divide-and-conquer algorithm yields slightly worse identification results but is computationally faster.
Figure 6: Community detection simulation results with Walktrap base algorithm. The number in the upper-left corner corresponds to the simulation setting (1-6).
Figure 7: Community detection simulation run-time results with Walktrap base algorithm. The number in the upper-left corner corresponds to the simulation setting (1-6).
Simulation sub-samples timing results
In Figures 8 and 9, we report the average time it takes to detect the community structure and CP structure on the sub-samples, respectively, for the main simulation settings. In general, RNN is the slowest with RN being the fastest.
Figure 8: Average computing time to detect the community structure for each sub-sample with Fast Greedy base algorithm. The number in the upper-left corner corresponds to the simulation setting (1-6).
Figure 9: Average computing time to detect the CP structure for each sub-sample. The number in the upper-left corner corresponds to the simulation setting (7-12).
Simulation sub-samples density results
We also report the average sub-graph density for each method across each simulation setting. For each setting and parameter combination, we generate 10 networks, and then sub-sample the network 100 times for each sub-sampling algorithm. We report the average density of the sub-sample, $\hat{p}^{(b)}$, where
$\hat{p}^{(b)}=\frac{1}{nq(nq-1)}\sum_{i,j}A^{(b)}_{ij}$
where ${{\bf A}}^{(b)}$ is the adjacency matrix for sub-sampled network $\mathcal{S}^{(b)}$. Note that settings 3 (9) and 5 (11) are redundant for the community (CP) simulations so they are excluded. We also plot the density of the original graph. The results are in Figures 10 and 11.
Figure 10: Average sub-graph density for community detection simulations. The number in the upper-left corner corresponds to the simulation setting (1-6).
Figure 11: Average sub-graph density for CP detection simulations. The number in the upper-left corner corresponds to the simulation setting (7-12).